OpenStudy (emmigrace222):

(5 + 2)^2 - 40 + 3 = __________. \( Numerical~Answers~Expected!\)

OpenStudy (anonymous):

(5+2)=7 7^2=49 49-40+3=12

OpenStudy (anonymous):

Use PEMDAS \[(5+2)^2-40+3\]\[7^2-40+3\]\[49-40+3\] which equals 12